Arthur and Betty start walking toward each other when they are 100 m apart. Arthur has a speed of 3.0 (2.0m / m / s and Betty has a speed of How long does it take for them to meet?

1
Expert's answer
2022-04-11T12:04:55-0400

"t=\frac s{v_1+v_2}=\frac{100}{3+2}=20~s."